# Hao Chen1, Bo Zhang1, Chunhua Shen1 1

### v1

@@ Appendix A Implementation Details
Appendix B Multi-modal Scripts Data Pipeline Details
In this chapter, we explain in detail how we construct the multimodal script data. Specifically, We collect a massive movie dataset, with most movies coming from Condensed Movies [1] and others from YouTube.
Figure 14: Multi-modal scripts data pipeline. We generate image captions using existing VLMs, and then further utilize LLMs to reformat the annotations into a structured script. We use face retrieval method to extract different face embeddings of the same character, which benefits anti-overfitting data augmentation in Section 3.3.
We first employ a keyframe extraction technique to sample frames from the video at the regular interval of 30 frames. Subsequently, we segment the top 5 characters with the highest confidence scores in each frame with Grounding-Dino [28] with prompt "person" and extract the face embedding of the segmented characters [48]. We then leverage an existing VLM [21] to generate captions for keyframes as well as the corresponding segmented characters. Lastly, we employ an LLM to fuse the captions of the keyframes and the corresponding characters to synthesize the comprehensive and structured multi-modal scripts, which corresponds to the Step 1 and Step 2 in Figure 15. For step 3, we employ a large language model to extract the scene elements and the plot of the corresponding frame. Finally, we reformat the character, scene elements and the plot into a structured representation. This pipeline effectively constructs large-scale multi-modal script dataset, providing a substantial amount of high-quality data for training. Script data provides rich contextual information, enabling the model to better remember and utilize this information to generate consistent content.
@@ Appendix B Multi-modal Scripts Data Pipeline Details
Appendix C Additional Results
In this section, we present additional generated stories and videos. Our method exhibits strong generalization capabilities, producing high-quality results with inputs not included in the training set. Our method can understand text and character inputs, capable of generating results that align with the storyline while ensuring character identity consistency. The overall scene consistency is also well-preserved, as our autoregressive model is able to context information as reference. Most importantly, we are the first method to generate extremely long contents while preserving both short-term and long-term consistency. Besides, we even ensure that the IDs of different characters remain consistent before and after a camera change.
Consistency Preservation.
Our method effectively preserves both short-term and long-term consistency for multiple characters. We present numerous results in Figure 9. Our method is able to preserve multiple characters within long contents without the need for any fine-tuning. Even when the scene changes, our method effectively maintains the same character ID before and after the shot transition.
@@ Appendix C Additional Results
We showcase our ability to preserve character ID in Figure 16. Our method exhibits strong generalization capabilities, allowing it to generate corresponding character IDs based on user input without any further training.
@@
Figure 16: ID Preservation results. Our method demonstrates great ability to preserve ID.
@@
The generated keyframes can serve as the condition to generate videos, as shown in Figure 17. Our keyframes have successfully preserved both short-term and long-term temporal consistency, facilitating the seamless rendering of these frames into a cohesive and harmonious video.
@@
Figure 17: Example video results.
Appendix D Additional Comparisons
@@
Appendix E Limitations and Future Works
E.1 Limitations
@@
Figure 18: When there are too many people in one frame, the model sometimes confuses the IDs.
@@
Despite our method demonstrating strong potential in generating long sequences, it still has some limitations. (1) We employ longCLIP to compress each sentence of descriptive texts into a single token. While this strategy effectively reduces the sequence length, it provides only a coarse representation of each sentence, which causes information loss. (2) Our method’s ability to preserve objects consistency is limited. We did not construct the multi-modal script with scene information, as the costs far exceed our capacity. Moreover, as shown in Fig. 18, if there are too many people in one frame, our method is prone to mixing the appearance of characters. (3) Analogous to other large models, our approach requires substantial data and computational power for training. (4) Currently, we generate 128 keyframes at a time and run multiple iterations of this process to create long contents. However, our goal is to extend the maximum token length, fundamentally addressing the long sequence generation by using long sequence generation methods in LLMs. Due to resource constraints, we are unable to pursue this approach at present and leave it in future work.
E.2 Future Works
Better sentence representation.
We call for a better sentence compression method which compresses sequence length while maximally preserving the semantic information of the original sentences. This would be particularly beneficial as it increases the number of visible frames.
Better keyframe-based video generation model.
We believe that the optimal solution for keyframe-to-video generation would be to design a model capable of conditioning on multiple keyframes. We have made some initial attempts, however, training such a model requires a substantial amount of data and computation overhead, which far exceeds the resources we can provide. We believe a model conditioning on multiple keyframes is vital for generating long videos and consider this as the future work.
Controllability beyond face identity.
Currently, our method primarily focuses on maintaining character consistency. Future work can further explore maintaining consistency in additional attributes such as objects and scenes. How to address the issue of character confusion is also an important area for futre works.
Long sequence generation.
We aim to fundamentally solve the problem of long sequence generation, rather than compromising with the current approach of iteratively generating 128 frames. Modern LLMs already support very long context tokens. We believe we can draw inspiration from these methods to achieve true generation of extremely long content.
E.3 Social impact
@@
Our method can generate high-quality long stories and videos, significantly lowering the barrier for individuals to create desired visual appealing content. However, it is essential to address the negative potential social impact of our method. Malicious users may exploit its capabilities to generate inappropriate or harmful content. Consequently, it is imperative to emphasize the responsible and ethical utilization of our method, under the collective supervision and governance of society as a whole. Ensuring appropriate usage practices requires a collaborative effort from various people, including researchers, policymakers, and the wider community.

### v2

@@ Appendix A Implementation Details
Appendix B Evaluation Metrics
We propose new metrics to evaluate short-term (ST) consistency and long-term (LT) consistency. For short-term consistency, we select all the frames where a character appears. We then calculate the similarity of CLIP embedding of selected consecutive keyframes, as consecutive keyframes where the same character appears usually have high similarity. For long-term consistency, We calculate the proportion of generated frames where the character appears to the ground truth frames where the character appears. For example, if the GT frames where a character appears are 1, 2, 3, 4, 5, and the generated frames where the character appears are 1, 2, 3, 6, then the long-term consistency is calculated as 3/5. To select the frames where a character appears, we first select an image of the main character and use FARL to obtain the character’s embedding. We then compute the similarity of this embedding with FARL embeddings of all other frames and select images above a certain threshold.
The Aesthetic Score (Schuhmann et al., 2022) is calculated based on the CLIP image features. Each image is first passed through the CLIP image encoder to obtain an embedding. This embedding is then fed into an additional small neural network, which predicts an aesthetic score.
Appendix C Multi-modal Scripts Data Pipeline Details
Multimodal script construction.
In this chapter, we explain in detail how we construct the multimodal script data. Specifically, We collect a massive movie dataset, with movies coming from Condensed Movies (Bain et al., 2020) and others collected using the same methodology as Condensed Movies. The data consists of around 23,000 long videos, which subsequently results in 5 million keyframes with corresponding script annotations. The overall multimodal script synthesis is as follows:
Figure 16: Multi-modal scripts data pipeline. We generate image captions using existing VLMs, and then further utilize LLMs to reformat the annotations into a structured script. We use the face retrieval method to extract different face embeddings of the same character, which benefits anti-overfitting data augmentation in Section 3.4.
  • Extract keyframes from long videos at 1-second intervals, removing the beginning and end segments. These segments primarily contain sponsor messages, acknowledgments, and advertisements, lacking useful information. Subsequently, center-crop and resize each frame to a width of 896 pixels and a height of 512 pixels. Finally, leverage a Vision Language Model (Lu et al., 2024) to generate descriptions for the entire image.
  • Using GroundingDINO (Ren et al., 2024a) with ”Person” as the prompt, extract the top 5 characters from each frame. If the number of characters exceeds 5, the extra characters are ignored. If the number of characters is less than 5, there is no additional processing.
  • Employ an existing Vision Language Model (Lu et al., 2024) to obtain captions for each segmented character. The face of each character is extracted to construct the face ID bank.
  • Use a Large Language Model, taking the captions for the entire image extracted in step 1 and captions for each character in step 3 as input, to derive Scene Elements and Plot, and ask the model to output in a multimodal script format. At this point, we have structured Scene Elements and Plot but still lack Character. We then format all the character captions corresponding to each frame into the Character section of the multimodal script.
  • However, the character face IDs for each frame are derived from the same frame, which can lead to the model overfitting by directly copy-pasting characters into the generated results. To address this, we use FARL (Zheng et al., 2022) to randomly retrieve the same character’s ID from the face ID bank as input. This strategy effectively mitigates the overfitting issue.
@@ Appendix C Multi-modal Scripts Data Pipeline Details
Appendix D Additional Results
In this section, we present additional generated stories and videos. Our method exhibits strong generalization capabilities, producing high-quality results with inputs not included in the training set. Our method can understand text and character inputs, capable of generating results that align with the storyline while ensuring character identity consistency. The overall scene consistency is also well-preserved, as our autoregressive model is able to leverage context information as the reference. Most importantly, we are the first method to generate extremely long contents while preserving both short-term and long-term consistency. Besides, we even ensure that the IDs of different characters remain consistent before and after a camera change.
Generation process.
Our autoregressive model is similar to a multi-turn dialogue model. Specifically, we input the multimodal script corresponding to a keyframe, and the model autoregressively outputs two compressed tokens. Then, based on this, we input the multimodal script for the next frame to obtain the next two compressed tokens, without losing historical information. After obtaining all the compressed tokens, we input them along with the face embeddings and description embeddings from the multimodal script into the decoder to generate the keyframe. The keyframes are subsequently leveraged by image-to-video model to generate videos.
Additional results.
Our method effectively preserves both short-term and long-term consistency for multiple characters. We present numerous results in Figure 10111213 and  14. Our method is able to preserve multiple characters within long contents without the need for any fine-tuning. Even when the scene changes, our method effectively maintains the same character ID even during camera switches. Due to space limitations, we only select certain excerpts to showcase here. Please refer to the TITANIC generated by our MovieDreamer in supplementary material for longer results.
@@ Appendix D Additional Results
We showcase our ability to preserve character ID in Figure 18. Our method exhibits strong generalization capabilities, allowing it to generate target characters without any further training.
@@
Figure 18: ID Preservation results. Our method demonstrates the great ability to preserve ID. The compressed token for each keyframe is generated with the corresponding multimodal script as input, and the keyframes are rendered with compressed tokens, face embedding and description embedding as input. We show the target ID on the left.
@@
The generated keyframes serve as the condition to generate videos. We show the results generated with our improved video generation model in Figure 19. Our keyframes have successfully preserved both short-term and long-term temporal consistency, facilitating the seamless rendering of these frames into a cohesive and harmonious video.
Again, we would like to highlight that our primary contribution lies in the hierarchical controllable very long generation of visual contents using MLLM for keyframe generation and diffusion model for video rendering. Any high-quality image-to-video model can be integrated with our approach. We showcase video results generated using higher-quality image-to-video models with our generated keyframes as input in Figure 20, which strongly demonstrates the effectiveness and flexibility of our method. We believe our approach is highly promising has immense potential.
Different shot types and styles.
We find that although we did not explicitly model the camera shot types, the model can implicitly learn various shot types based on the multimodal script. The model is capable of selecting the most appropriate shot type based on the input. The shot types of anchor frames further assist the image-to-video model in generating videos with various shot types. Our method is also capable of generating cartoon results. We showcase the results in Figure 21.
@@
Figure 19: Example video results with our video generation method.
Figure 20: Example video results using existing better image-to-video model LUMA. Existing image-to-video models can be integrated with our method to produce high-quality ultra-long videos, which demonstrates the flexibility and effectiveness of our MovieDreamer.
Figure 21: (a) Although our method does not explicitly model shot types, it can still implicitly learn various shot types and generate high-quality results with suitable shot types, demonstrating the potential of our approach. (b) Our method can also generate cartoon results and interleaved characters.
Appendix E Additional Experiments
E.1 Additional Comparisons
@@
E.2 Additional Ablations on the Number of Compressed Tokens
Our goal and an important tradeoff.
While more tokens may yield better results, they sacrifice the context length of autoregression. Since we aim for long video generation, using as few tokens as possible is a reasonable design choice. Our goal is to achieve the highest quality images with a minimal number of tokens.
Why do we choose 2 tokens.
Existing work Paint-by-Example (Yang et al., 2022) demonstrates that a single token can effectively reconstruct the original image. Building on this discovery, we added a few additional tokens. We compared the reconstructed images using 8 tokens and 2 tokens, with the metrics shown in the accompanying table below. The metrics indicate that increasing the number of tokens does not offer significant improvements in results. As shown in Table 3, we calculate the inception score, aesthetic score as well as FID score of images of different numbers of tokens. We further compare the CLIP similarity between reconstructed images and target images. The evaluation metrics demonstrate that the results of 8 tokens are slightly improved. But considering our objective of long content generation, this slight improvement is not enough for us to prioritize it when making trade-offs regarding sequence length and our limited computational resources, as 8 tokens will lead to a much shorter sequence length and demand more computational overhead. Consequently, we eventually chose 2 tokens. But we believe that if there are enough computational resources, 8 tokens will be a better choice.
Table 3: Quantitative ablation study on the number of compressed tokens.
[…]
2 tokens 0.825 8.235 6.378 0.406
8 tokens 0.841 8.193 6.457 0.348
Appendix F Limitations and Future Works
@@
Figure 22: When there are too many characters in one frame, the model sometimes confuses the character IDs.
@@
Despite our method demonstrating strong potential in generating long sequences, it still has some limitations. (1) We employ LongCLIP to compress each sentence of descriptive text into a single token. While this strategy effectively reduces the sequence length, it provides only a coarse representation of each sentence, which causes information loss. (2) Our method’s ability to preserve object consistency is limited. We did not construct the multi-modal script with scene information, as the costs far exceed our capacity. Moreover, as shown in Fig. 22, if there are too many people in one frame, our method is prone to mixing the appearance of characters. (3) Analogous to other large models, our approach requires substantial data and computational power for training. (4) Currently, we generate 128 keyframes at a time and run multiple iterations of this process to create long content. However, our goal is to extend the maximum token length, fundamentally addressing the long sequence generation by using long sequence generation methods in LLMs. Due to resource constraints, we are unable to pursue this approach at present and leave it in future work.
Social impact.
@@
Our method can generate high-quality long stories and videos, significantly lowering the barrier for individuals to create desired visually appealing content. However, it is essential to address the negative potential social impact of our method. Malicious users may exploit its capabilities to generate inappropriate or harmful content. Consequently, it is imperative to emphasize the responsible and ethical utilization of our method, under the collective supervision and governance of society as a whole. Ensuring appropriate usage practices requires a collaborative effort from various people, including researchers, policymakers, and the wider community. Furthermore, the code, model, as well as the data, will be fully released to improve the development of related fields.